OpenStudy (sophadof):

Solve by completing the square: x^2+6x=12

OpenStudy (anonymous):

add both sides (6/2)^2 i.e., 9 and complete

OpenStudy (jdoe0001):

did you get what surjithayer suggested?

OpenStudy (jdoe0001):

yes? no? maybe? no the foggiest? sorry was busy watching spongebob? just had an ice-cream headache?

OpenStudy (sophadof):

Uh... not really..

OpenStudy (danjs):

x^2 + 2ax + a^2 = (x + a)^2 that is what you want to do, make it so you can do that. you have x^2 + 6x + ... 2a = 6 in this case to get the 'a^2' value to add on, you have to take half of that 6, a=3, then square that a^2=9

OpenStudy (danjs):

x^2 + 6x + 9 -----> (x+3)^2 make sure you also add 9 to the other side, to not screw up the equality

OpenStudy (danjs):

The overall thing to remember in the case where you have x^2 + ax take half of that 'a' value, and square it, add it to both sides of equation...

OpenStudy (danjs):

try complete the square x^2 + 20x = 7

OpenStudy (sophadof):

I uh.. am still trying to do this one hehe @DanJS

OpenStudy (sophadof):

I think I got the first one!! Is it.. x^2+6x+9=x+9 ?

OpenStudy (danjs):

x^2 + 6x =12 half the a term and square it... 6/2 = 3 , 3^2 = 9, add to both sides x^2 + 6x + 9 = 12 + 9 makes it able to be a square (x + 3)(x+3) = (x+3)^2 = 21

OpenStudy (danjs):

yep

OpenStudy (sophadof):

YAY! thank you so much! That helped a lot!! I will keep this in my notes... @DanJS

OpenStudy (danjs):

what is x^2 + 20x = 7 as a coimpleted square?